# LaDe: The First Comprehensive Last-mile Delivery Dataset from Industry

### v1

@@ 4.1 Route Prediction
Table 3: Experimental results of route prediction. We use bold and underlined fonts to denote the best and runner-up model, respectively.
@@
HR@3 $\uparrow$ KRC $\uparrow$ LSD $\downarrow$ ED $\downarrow$ HR@3 $\uparrow$ KRC $\uparrow$ LSD $\downarrow$ ED $\downarrow$ HR@3 $\uparrow$ KRC $\uparrow$ LSD $\downarrow$ ED $\downarrow$
TimeGreedy 63.86 44.16 3.91 1.74 57.65 31.81 5.54 2.15 61.23 39.64 4.62 1.85
DistanceGreedy 62.99 41.48 4.22 1.60 60.77 39.81 5.54 2.15 62.34 40.82 4.49 1.64
OR-Tools 64.19 43.09 3.67 1.55 66.21 47.60 4.40 1.81 63.27 42.31 3.94 1.59
LightGBM 71.55 54.53 2.63 1.54 73.76 55.71 3.01 1.84 70.41 52.90 2.87 1.59
FDNET 69.98 $\pm$ 0.32 52.07 $\pm$ 0.38 3.36 $\pm$ 0.04 1.51 $\pm$ 0.01 73.27 $\pm$ 0.47 53.80 $\pm$ 0.58 3.30 $\pm$ 0.04 1.84 $\pm$ 0.01 69.08 $\pm$ 0.61 50.62 $\pm$ 1.20 3.60 $\pm$ 0.15 1.57 $\pm$ 0.02
DeepRoute 72.09 $\pm$ 0.39 55.72 $\pm$ 0.40 2.66 $\pm$ 0.06 1.51 $\pm$ 0.01 74.68 $\pm$ 0.07 56.60 $\pm$ 0.16 2.98 $\pm$ 0.01 1.79 $\pm$ 0.01 71.44 $\pm$ 0.28 54.74 $\pm$ 0.49 2.80 $\pm$ 0.02 1.53 $\pm$ 0.02
Graph2Route 72.31 $\pm$ 0.20 56.08 $\pm$ 0.14 2.53 $\pm$ 0.01 1.50 $\pm$ 0.01 74.84 $\pm$ 0.15 56.99 $\pm$ 0.52 2.86 $\pm$ 0.02 1.77 $\pm$ 0.01 71.52 $\pm$ 0.14 55.02 $\pm$ 0.10 2.71 $\pm$ 0.01 1.54 $\pm$ 0.01
@@ 4.2 Estimated Time of Arrival Prediction
Results. MAE and RMSE, and ACC@30 are used to evaluate the performance of time prediction models. Higher ACC@30 and lower MAE and RMSE indicate better performance. From the results shown in Table 4, we can see that deep models outperform the shallow models by a large margin, because of their ability to model multiple spatio-temporal factors. We also observe a huge performance gap of the same method in different cities. For example, the best model, FDNET, achieves 84% in terms of ACC@30 in Shanghai while it gets a much lower accuracy of 66% in another two datasets. It deserves further study to build a more powerful model that can generalize well in cities with different properties.
@@
Method Shanghai Chongqing Yantai
MAE $\downarrow$ RMSE $\downarrow$ ACC@30 $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ ACC@30 $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ ACC@30 $\uparrow$
LightGBM 30.99 35.04 0.59 43.09 48.93 0.44 44.19 50.20 0.42
SPEED 23.75 27.86 0.73 33.42 39.45 0.58 31.41 37.09 0.60
KNN 36.00 31.89 0.58 51.55 45.68 0.40 51.55 45.68 0.40
MLP 21.54 $\pm$ 2.2 25.05 $\pm$ 2.46 0.79 $\pm$ 0.04 32.38 $\pm$ 0.92 37.75 $\pm$ 1.25 0.66 $\pm$ 0.01 27.97 $\pm$ 0.97 32.35 $\pm$ 1.16 0.65 $\pm$ 0.01
FDNET 18.47 $\pm$ 0.25 21.44 $\pm$ 0.28 0.84 $\pm$ 0.01 28.17 $\pm$ 1.16 32.60 $\pm$ 1.36 0.65 $\pm$ 0.01 27.08 $\pm$ 2.8 31.15 $\pm$ 3.47 0.65 $\pm$ 0.01

### v3

@@ 4.1 Route Prediction
Table 3: Experiment Results of Route Prediction.
@@
HR@3 KRC LSD ED HR@3 KRC LSD ED HR@3 KRC LSD ED
Time-Greedy 63.86 $\pm$0.00 44.16 $\pm$0.00 3.91 $\pm$0.00 1.74 $\pm$0.00 59.81 $\pm$0.00 39.93 $\pm$0.00 5.20 $\pm$0.00 2.24 $\pm$0.00 61.23 $\pm$0.00 39.64 $\pm$0.00 4.62 $\pm$0.00 1.85 $\pm$0.00
Distance-Greedy 62.99 $\pm$0.00 41.48 $\pm$0.00 4.22 $\pm$0.00 1.60 $\pm$0.00 61.07 $\pm$0.00 42.84 $\pm$0.00 5.35 $\pm$0.00 1.94 $\pm$0.00 62.34 $\pm$0.00 40.82 $\pm$0.00 4.49 $\pm$0.00 1.64 $\pm$0.00
Or-Tools 64.19 $\pm$0.00 43.09 $\pm$0.00 3.67 $\pm$0.00 1.55 $\pm$0.00 62.50 $\pm$0.00 44.81 $\pm$0.00 4.69 $\pm$0.00 1.88 $\pm$0.00 63.27 $\pm$0.00 42.31 $\pm$0.00 3.94 $\pm$0.00 1.59 $\pm$0.00
LightGBM 71.55 $\pm$0.00 54.53 $\pm$0.00 2.63 $\pm$0.00 1.54 $\pm$0.00 70.63 $\pm$0.00 54.48 $\pm$0.00 3.27 $\pm$0.00 1.92 $\pm$0.00 70.41 $\pm$0.00 52.90 $\pm$0.00 2.87 $\pm$0.00 1.59 $\pm$0.00
FDNET 69.98 $\pm$0.32 52.07 $\pm$0.38 3.36 $\pm$0.04 1.51 $\pm$0.01 69.05 $\pm$1.23 52.72 $\pm$1.72 4.08 $\pm$0.25 1.86 $\pm$0.03 69.08 $\pm$0.61 50.62 $\pm$1.20 3.60 $\pm$0.15 1.57 $\pm$0.02
DeepRoute 72.09 $\pm$0.39 55.72 $\pm$0.40 2.66 $\pm$0.06 1.51 $\pm$0.01 71.66 $\pm$0.10 56.20 $\pm$0.23 3.26 $\pm$0.07 1.86 $\pm$0.01 71.44 $\pm$0.28 54.74 $\pm$0.49 2.80 $\pm$0.02 1.53 $\pm$0.02
CPRoute 72.55 $\pm$0.23 55.76 $\pm$0.32 2.70 $\pm$0.01 1.49 $\pm$0.02 71.73 $\pm$0.08 56.17 $\pm$0.04 3.39 $\pm$0.02 1.84 $\pm$0.00 71.76 $\pm$0.04 54.84 $\pm$0.03 2.99 $\pm$0.01 1.53 $\pm$0.00
Graph2Route 72.31 $\pm$0.20 56.08 $\pm$0.14 2.53 $\pm$0.01 1.50 $\pm$0.01 71.69 $\pm$0.10 56.53 $\pm$0.10 3.12 $\pm$0.01 1.86 $\pm$0.00 71.52 $\pm$0.14 55.02 $\pm$0.10 2.71 $\pm$0.01 1.54 $\pm$0.00
M2G4RTP 72.44 $\pm$0.11 56.15 $\pm$0.04 2.55 $\pm$0.03 1.50 $\pm$0.00 71.73 $\pm$0.06 56.34 $\pm$0.08 3.16 $\pm$0.02 1.86 $\pm$0.01 71.73 $\pm$0.01 55.02 $\pm$0.17 2.78 $\pm$0.00 1.53 $\pm$0.01
DRL4Route 73.12 $\pm$0.06 57.23 $\pm$0.12 2.43 $\pm$0.01 1.48 $\pm$0.01 72.18 $\pm$0.15 57.20 $\pm$0.18 3.06 $\pm$0.02 1.84 $\pm$0.01 72.07 $\pm$0.06 55.94 $\pm$0.10 2.62 $\pm$0.00 1.51 $\pm$0.00
@@ 4.2 Estimated Time of Arrival Prediction
[…]
@@
Method Shanghai Yantai Chongqing
MAE $\downarrow$ RMSE $\downarrow$ ACC@20 $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ ACC@20 $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ ACC@20 $\uparrow$
SPEED 26.68 $\pm$0.00 31.31 $\pm$0.00 52.57 $\pm$0.00 33.97 $\pm$0.00 40.27 $\pm$0.00 42.03 $\pm$0.00 35.55 $\pm$0.00 42.06 $\pm$0.00 41.10 $\pm$0.00
KNN 25.22 $\pm$0.00 29.57 $\pm$0.00 65.71 $\pm$0.00 28.10 $\pm$0.00 33.80 $\pm$0.00 45.33 $\pm$0.00 29.45 $\pm$0.00 35.19 $\pm$0.00 43.68 $\pm$0.00
LightGBM 17.24 $\pm$0.00 20.40 $\pm$0.00 67.44 $\pm$0.00 23.32 $\pm$0.00 27.82 $\pm$0.00 51.22 $\pm$0.00 24.22 $\pm$0.00 27.99 $\pm$0.00 48.80 $\pm$0.00
MLP 16.16 $\pm$0.02 19.31 $\pm$0.01 72.17 $\pm$0.10 22.18 $\pm$0.06 26.38 $\pm$0.06 54.98 $\pm$0.17 23.82 $\pm$0.02 28.24 $\pm$0.02 52.74 $\pm$0.16
FDNET 18.81 $\pm$1.23 21.15 $\pm$2.47 64.30 $\pm$1.43 22.41 $\pm$0.50 26.00 $\pm$0.63 54.67 $\pm$1.55 22.54 $\pm$0.91 24.53 $\pm$0.92 46.65 $\pm$4.82
RankETPA 15.76 $\pm$0.05 19.13 $\pm$0.08 73.67 $\pm$0.15 21.36 $\pm$0.07 25.52 $\pm$0.08 57.83 $\pm$0.07 23.37 $\pm$0.01 27.93 $\pm$0.04 54.45 $\pm$0.10
M2G4RTP 8.23 $\pm$0.02 9.59 $\pm$0.03 91.01 $\pm$0.51 17.21 $\pm$1.57 19.71 $\pm$1.88 69.20 $\pm$0.39 17.96 $\pm$0.25 19.74 $\pm$0.31 65.97 $\pm$0.32
@@ A.1 Data Field
Table 9: Description of data fields of Detailed Trajectory.
Data field Description Unit/format
ds The date of the trajectory Date
courier_id Id of the courier Id
gps_time The time when the trajectory point is recorded Time
lng/lat Coordinates of the courier Float
Table 10: Description of data fields of Road Network.
Data field Description Unit/format
id Unique identifier of the road Id
road_id Id of a street or place name Id
code 4 digit code (between 1000 and $9$) defining the geographical features of a road Id
fclass Class name of the road String
ref Reference number of the road String
oneway Whether the road is a oneway road String
maxspeed Max allowed speed in km/h Int
layer Relative layering of roads Int
bridge Whether the road is on a bridge String
tunnel Whether the road is in a tunnel String
city City of the road String
geometry A geometry type composed of one or more line segments String
@@ A.2 Data Statistics
Table 13: Statistics of Detailed Trajectory.
City Time span Spatial span #GPS points #Couriers
Shanghai 1 month 77km×$\times$73km 9727k 245
Hangzhou 1 months 211km×$\times$142km 17900k 443
Chongqing 1 months 403km×$\times$267km 13520k 658
Yantai 1 months 155km×$\times$114km 6616k 245
Jilin 1 months 105km×$\times$165km 2849k 613